OpenStudy (anonymous):

Use implicit differentiation to find dy/dx. ln xy + 5x = 30. Please write out the steps.

hartnn (hartnn):

you need to use chain rule for ln(xy) and then product rule for xy. try it. if you can't get, i'll help you further.

OpenStudy (anonymous):

d/dx (ln xy +5x) =d/dx (30) d/dx (ln xy) + d/dx (5x) =0 d/d(xy) (lnxy) * d/dx(xy) +5=0 1/xy * { d/dx(x) *y + x * d/dx(y)}=-5 y+x *dy/dx =-5xy x*dy/dx =-5xy-y dy/dx = -(5xy+y)/x